Last year ABC Corp had sales of $315,000, net income of $12,600,year end assets of $210,000. The firm's total liabilities to total assets ratio is 20%. Based on the DuPont equation what is the ROE?
Answer #1

Profit margin=Net income/Sales

=(12600/315,000)

=4%

Total liabilities to total assets=Total liabilities/total assets

Total liabilities=(210,000*20%)=$42000

total assets=Total liabilities+Total equity

Total equity=(210,000-42,000)=$168,000

Total asset turnover=Sales/Total assets

=(315,000/210,000)

=1.5

Equity multiplier=Total assets/Equity

=210,000/168,000

=1.25

ROE=Profit margin*Total asset turnover*Equity multiplier

=4%*1.5*1.25

=7.5%